40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 7 23:06:30 2018, Koji, Update, PSL, EOM measuement preparation 
    Reply  Wed Aug 8 20:56:11 2018, Koji, Update, PSL, EOM measuement preparation 
Message ID: 14145     Entry time: Wed Aug 8 20:56:11 2018     In reply to: 14144
Author: Koji 
Type: Update 
Category: PSL 
Subject: EOM measuement preparation 

Rich and I worked on the EOM measurement. After the measurement, the setup was reverted to the nominal state

  • AUX PLL mixer was restored to ZAD-6
  • The PLL gain was restored to 3.10
  • The main PSL marconi is connected to the freq generator again. Using the beat note, I've confirmed that the modulations are applied on the beam.
  • The PSL HEPA was reduced from 100 to 30.
ELOG V3.1.3-